OpenStudy (anonymous):

PLEASE HELP!!!!!! Which of the following exponential functions goes through the points (1, 12) and (2, 36)? f(x) = 4(3)−x f(x) = 3(4)−x f(x) = 3(4)x f(x) = 4(3)x I think its either C or D

OpenStudy (misty1212):

it is either C or D

OpenStudy (misty1212):

both give \(12\) if \(x=1\) so try \(x=2\) and see which one gives \(36\)

OpenStudy (misty1212):

which one equals 36? \[f(2)=3\times 4^2\] or \[f(2)=4\times 3^2\]?

OpenStudy (anonymous):

Hold on im going to work it out

OpenStudy (anonymous):

f(2)=4x9 is the one that gives me 36

OpenStudy (misty1212):

yeah pick D

OpenStudy (anonymous):

f(2)=4x3^2 is the one sorry lol

OpenStudy (misty1212):

you got it

OpenStudy (anonymous):

Thanks !:)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

Gave you a medal :D

OpenStudy (misty1212):

gotta run to class ttyl